OpenStudy (anonymous):

1/x+2/y-4/z=1 2/x+3/y+8/z=0 -1/x+9/y+10/z=5 solve the system for x,y,z

OpenStudy (anonymous):

\[\Large \frac1x+\frac2y-\frac4z=1\]\[\Large \frac2x+\frac3y+\frac8z=0\]\[\Large-\frac1x+\frac9y+\frac{10}z=5\]

OpenStudy (anonymous):

This really hurts to look at ^.^ How about you do this, instead? \[\Large x^{-1}+2y^{-1}-4z^{-1}=1\]\[\Large 2x^{-1}+3y^{-1}+8z^{-1}=0\]\[\Large -x^{-1}+9y^{-1}+10z^{-1}=5\]

OpenStudy (saifoo.khan):

:O

OpenStudy (anonymous):

And now it's down to bread-and-butter systems :) There are sooo many ways to do this, but since you said Linear Algebra, what say we use Kramer's Rule? Think of adventure.......^.^

OpenStudy (anonymous):

pete pan now i got idea thanks

OpenStudy (anonymous):

That's music to my ears... even though I didn't hear a thing ^.^

OpenStudy (aravindg):

great :) ....you can also use matrix method if you wish !

OpenStudy (saifoo.khan):

How about this? \[\frac x1 + \frac y2 - \frac z4 = \frac 11\]\[\frac x2 + \frac y3 - \frac z8 = 0\]\[-\frac x1 + \frac y9 - \frac {z}{10}= \frac 15\] 1/x+2/y-4/z=1 2/x+3/y+8/z=0 -1/x+9/y+10/z=5

OpenStudy (anonymous):

but its too complicated

OpenStudy (anonymous):

Hey, @saifoo.khan That really hurts :D You're not suggesting \[\huge a+b+c=d\] means \[\huge a^{-1}+b^{-1}+c^{-1}=d^{-1}\] are you...? naughty...

OpenStudy (aravindg):

yaa it hurts badly :/

OpenStudy (anonymous):

For which values of a will the followingsystem have no solutions? Exactly one solution? Infinitely many solutions? x + 2y − 3z = 4 3x − y + 5z = 2 4x + y + (a2−14)z = a +2

OpenStudy (anonymous):

Interesting...

OpenStudy (anonymous):

interseting why?

OpenStudy (anonymous):

Nothing... I'm trying to see if I'm correct :)

OpenStudy (anonymous):

Okay, I know the answer. That was fun ^.^

OpenStudy (anonymous):

but i dnt know the answer:(

OpenStudy (anonymous):

This was played real nicely :)

OpenStudy (anonymous):

So, anyway, I want to know if you have an idea, first :)

OpenStudy (anonymous):

ya i know the idea of reduce echwelon form

OpenStudy (anonymous):

Oh. Is that so? What happens when you add the first two equations?

OpenStudy (anonymous):

ya its like reduce row echelon form

OpenStudy (anonymous):

Answer me....>.> What's the sum of the first two equations?

OpenStudy (anonymous):

1 2 -3 4 0 1 -2 10 0 0 1 a^2-4/a^2-16 i have done till this but now the problem is that For which values of a will the followingsystem have no solutions? Exactly one solution? Infinitely many solutions? how to find this

OpenStudy (anonymous):

\[\huge \left[ \begin{array}{ccc} 1&2&-3\\ 3&-1&5\\ 4&1&a^2-14 \end{array} \right| \left. \begin{array}{ccc} 4\\ 2\\ a+2 \end{array} \right]\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You could subtract the first and second row from the third.

OpenStudy (anonymous):

i know that we want a solution have you read that question

OpenStudy (anonymous):

I did :)

OpenStudy (anonymous):

so then

OpenStudy (anonymous):

So then... subtract the first and the second row from the third ^.^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok? What did you get?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

Well, you better get something :D

OpenStudy (anonymous):

Or, you can forget the matrix, and just add the first two equations ^.^

OpenStudy (anonymous):

hey you dont get my point i know all that but u dnt understand question

OpenStudy (anonymous):

What don't I get?